### v2

# Exact Factorization of Unitary Transformations with Spin-Adapted Generators

Abstract
Preserving spin symmetry in variational quantum algorithms is essential for producing physically meaningful electronic wavefunctions. Implementing spin-adapted transformations on quantum hardware, however, is challenging because the corresponding fermionic generators translate into noncommuting Pauli operators. In this work, we introduce an exact and computationally efficient factorization of spin-adapted unitaries derived from fermionic double excitation and deexcitation rotations. These unitaries are expressed as ordered products of exponentials of Pauli operators. Our method exploits the fact that the elementary operators in these generators form small Lie algebras. By working in the adjoint representation of these algebras, we reformulate the factorization problem as a low-dimensional nonlinear optimization over matrix exponentials. This approach enables precise numerical reparametrization of the unitaries without relying on symbolic manipulations. The proposed factorization provides a practical strategy for constructing symmetry-conserving quantum circuits within variational algorithms. It preserves spin symmetry by design, reduces implementation cost, and ensures the accurate representation of electronic states in quantum simulations of molecular systems.
preprint: APS/123-QED
I Introduction
Preserving spin and particle-number symmetries in variational quantum algorithms (VQAs) is essential for accurately describing molecular electronic structure. These symmetries ensure physically consistent potential energy surfaces, accelerate convergence, and prevent artificial symmetry breaking. However, maintaining them within scalable ansätze remains a significant challenge for quantum simulations of electronic systems.
Conventional unitary coupled cluster (UCC) ansätze[1] and their adaptive extensions[13, 11, 5, 14, 3] construct the variational wavefunction from a pool of anti-Hermitian generators chosen to balance expressiveness and circuit depth. In principle, the complete set of $4^{N}-1$ traceless Pauli generators acting on $N$ qubits spans the full $\mathfrak{su}(2^{N})$ Lie algebra, guaranteeing convergence to the exact ground state within a fixed orbital basis. The exponential scaling of this set, however, makes its direct use computationally infeasible. To achieve polynomial scaling, smaller operator pools that are closed under commutation have been proposed in both fermionic[5] and qubit operator spaces[14]. While such reduced pools can in principle generate the full algebra, many of them break key physical symmetries, including total spin ($S^{2}$), its projection ($S_{z}$), and particle number ($N_{e}$).
Maintaining these symmetries during variational optimization is desirable not only for physical consistency but also for improved convergence and reduced redundancy in the variational parameters. Several strategies have been proposed to enforce or restore symmetry, including penalty-based constraints[12], projection methods[17], and modified symmetry-conserving generators[2, 3]. The DISCO framework[3], for example, introduces one-electron and electron-pair rotations that preserve total spin by construction. Although effective for spin conservation, this approach restricts the operator pool to zero-seniority excitations (pair rotations), thereby reducing the overall expressivity of the ansatz.
Achieving full spin conservation while retaining the expressivity associated with all symmetrized double excitation and deexcitation rotations remains an open challenge. Implementing such transformations on quantum hardware requires their factorization into sequences of unitaries that can be compiled into standard gate sets. In this work, we address this gap by developing a numerically efficient and exact factorization scheme for unitaries generated by spin-adapted double excitation and deexcitation operators.
The primary difficulty in implementing these spin-adapted transformations arises because the corresponding generators are linear combinations of noncommuting operators in both fermionic and qubit representations. Standard Trotter-based decompositions therefore produce only approximate unitaries. The key insight of this work is that double excitation and deexcitation fermionic operators form compact Lie algebras. By exploiting the Lie algebraic closure, we rewrite the exponential of a sum of elementary fermionic generators as an ordered product of exponentials of the algebra’s basis elements, each of which can be implemented directly as a quantum circuit.
Since the number of basis elements in the relevant Lie algebras is modest (fewer than 100), the entire factorization can be carried out efficiently in the adjoint representation using low-dimensional matrix algebra. Instead of employing symbolic Baker–Campbell–Hausdorff expansions, we reduce the reparametrization problem to a small nonlinear optimization involving matrix exponentials in the adjoint representation. This numerical treatment achieves high precision while avoiding symbolic complexity.
The remainder of this paper presents the theoretical framework for this reparametrization, analyzes the structure of the relevant Lie algebras for different seniority sectors, and demonstrates numerical factorization for representative spin-adapted double excitation and deexcitation transformations. We conclude by discussing the implications of this factorization for constructing compact, symmetry-preserving variational ansätze in quantum chemistry.
II Method
II.1 Lie algebra framework
The goal of this work is to express unitaries that cannot be directly compiled into gate sequences as products of unitaries whose gate decompositions are known. Unitaries with spin-symmetrized generators can be written as
$\exp{\left[\sum_{k}c_{k}A_{k}\right]},$
where $A_{k}\in\mathcal{A}$ form a Lie algebra generated by the commutator closure of the $\{A_{k}\}$ operators. The dimensionality of these algebras, $m\lesssim 84$, depends on the seniority and the specific index pattern. Each separate exponential $\exp(tA_{k})$ admits a well-known gate representation, and therefore our objective is to decompose
$e^{\theta(\sum_{k}c_{k}A_{k})}$ $=\prod_{p}e^{t_{p}(\theta)A_{p}}.$ (1)
The first question is whether such a product representation exists for the Lie algebras considered. This issue is nontrivial because certain Lie algebras contain generator sets for which the exponential of a linear combination cannot be represented as a finite product of exponentials of the individual generators [4]. In our case, all generators are anti-Hermitian operators that form subalgebras of the compact $\mathfrak{su}(N)$ algebra, for which the exponential map is surjective. Consequently, any element of the Lie group generated by the left-hand side (LHS) of Eq. (1) can be expressed by an appropriate choice of parameters in the right-hand side (RHS).
To perform the decomposition, we employ a low-dimensional faithful matrix representation of the Lie algebras rather than symbolic Baker–Campbell–Hausdorff (BCH) expansions. This converts the factorization into a numerical matrix optimization problem, solved by minimizing the norm of the difference between the LHS and RHS of Eq. (1). A well-known alternative is the Wei–Norman reparametrization approach, which leads to a system of differential equations whose number equals the dimensionality of the Lie algebra. In our case, the algebras are large enough that analytical solutions are impractical, making a numerical approach more efficient. Moreover, the Wei–Norman method requires continuous trajectories $t_{p}(\theta)$ from $\theta=0$ to the target value, which may impose additional constraints on the Lie algebra beyond the surjectivity of the exponential map.
II.2 Lie algebras
We consider three Lie algebras corresponding to the following cases: seniority-2, seniority-4 intermediate singlet, and seniority-4 intermediate triplet. These algebras arise from the commutator closures of two-electron rotation operators $\hat{G}_{i_{\sigma}j_{\tau}}^{a_{\mu}b_{\nu}}$ present in every symmetrized generator:
$\hat{G}_{i_{\sigma}j_{\tau}}^{a_{\mu}b_{\nu}}=\hat{a}_{a_{\mu}}^{\dagger}\hat{a}_{b_{\nu}}^{\dagger}\hat{a}_{j_{\tau}}\hat{a}_{i_{\sigma}}-\hat{a}_{i_{\sigma}}^{\dagger}\hat{a}_{j_{\tau}}^{\dagger}\hat{a}_{b_{\nu}}\hat{a}_{a_{\mu}},$ (2)
where $i,j,a,b$ are spatial orbital indices and $\sigma,\tau,\mu,\nu$ are spin projection indices. The corresponding unitary transformations are generated by symmetrized combinations of double excitations and deexcitations.
Seniority-2:
$\hat{G}^{\mathrm{SA}}_{iiab}$ $=$ $\frac{1}{\sqrt{2}}\Big(\hat{G}^{a_{\alpha}b_{\beta}}_{i_{\alpha}i_{\beta}}-\hat{G}^{a_{\beta}b_{\alpha}}_{i_{\alpha}i_{\beta}}\Big),$ (3)
$\hat{G}^{\mathrm{SA}}_{ijaa}$ $=$ $\frac{1}{\sqrt{2}}\Big(\hat{G}^{a_{\alpha}a_{\beta}}_{i_{\alpha}j_{\beta}}-\hat{G}^{a_{\alpha}a_{\beta}}_{i_{\beta}j_{\alpha}}\Big).$ (4)
Seniority-4, intermediate singlet:
$\hat{G}^{\mathrm{SA}}_{ijab}=\frac{1}{2}\Big(\hat{G}^{a_{\alpha}b_{\beta}}_{i_{\alpha}j_{\beta}}+\hat{G}^{a_{\beta}b_{\alpha}}_{i_{\beta}j_{\alpha}}-\hat{G}^{a_{\alpha}b_{\beta}}_{i_{\beta}j_{\alpha}}-\hat{G}^{a_{\beta}b_{\alpha}}_{i_{\alpha}j_{\beta}}\Big),$ (5)
Seniority-4, intermediate triplet:
$\hat{G}^{\mathrm{SA}^{\prime}}_{ijab}$ $=\frac{1}{2\sqrt{3}}\Big(\hat{G}^{a_{\alpha}b_{\beta}}_{i_{\alpha}j_{\beta}}+\hat{G}^{a_{\beta}b_{\alpha}}_{i_{\beta}j_{\alpha}}+\hat{G}^{a_{\alpha}b_{\beta}}_{i_{\beta}j_{\alpha}}+\hat{G}^{a_{\beta}b_{\alpha}}_{i_{\alpha}j_{\beta}}$
$\qquad\qquad+2\,\hat{G}^{a_{\alpha}b_{\alpha}}_{i_{\alpha}j_{\alpha}}+2\,\hat{G}^{a_{\beta}b_{\beta}}_{i_{\beta}j_{\beta}}\Big).$ (6)
We summarize the closures for all three cases, providing explicit expressions for the seniority-2 algebra in this section and the seniority-4 algebras in Appendices AB. All algebras share several structural features, which we illustrate using the seniority-2 example.
First, the commutator closure of generators of the form $\hat{G}_{i_{\sigma}j_{\tau}}^{a_{\mu}b_{\nu}}$ produces additional elements of the type $p[\{\hat{n}(q_{\sigma})\}]\hat{G}_{i_{\sigma}j_{\tau}}^{a_{\mu}b_{\nu}}$, where $p[\{\hat{n}(q_{\sigma})\}]$ are polynomials of occupation number operators $\hat{n}(p_{\sigma})=\hat{a}^{\dagger}_{p_{\sigma}}\hat{a}_{p_{\sigma}}$. These polynomials have only three eigenvalues $\{\pm 1,0\}$. Their indices do not overlap with those in $\hat{G}_{i_{\sigma}j_{\tau}}^{a_{\mu}b_{\nu}}$, and the two factors in each element commute, $p[\{\hat{n}(q_{\sigma})\}],\hat{G}_{i_{\sigma}j_{\tau}}^{a_{\mu}b_{\nu}}]=0.$ Because both factors have spectra $\{\pm 1,0\}$, the total spectrum of each Lie algebra generator is also $\{\pm 1,0\}$.
For example, one basis of the seniority-2 Lie algebra is
$A_{1}$ $=\hat{G}^{a_{\alpha}b_{\beta}}_{i_{\alpha}i_{\beta}},$ (7)
$A_{2}$ $=\hat{G}^{a_{\beta}b_{\alpha}}_{i_{\alpha}i_{\beta}},$ (8)
$A_{3}$ $=(1-\hat{n}(i_{\alpha})-\hat{n}(i_{\beta}))\,\hat{G}^{a_{\beta}b_{\alpha}}_{a_{\alpha}b_{\beta}},$ (9)
$A_{4}$ $=-\Big(1-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})+2\hat{n}(a_{\alpha})\hat{n}(b_{\beta})\Big)\,\hat{G}^{a_{\beta}b_{\alpha}}_{i_{\alpha}i_{\beta}},$ (10)
$A_{5}$ $=\Big(1-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})+2\hat{n}(a_{\beta})\hat{n}(b_{\alpha})\Big)\,\hat{G}^{a_{\alpha}b_{\beta}}_{i_{\alpha}i_{\beta}},$ (11)
such that $\mathcal{A}_{S2}=\{A_{i}\}_{i=1}^{5}$. These basis elements satisfy the commutation relations
$[A_{1},A_{2}]=A_{3},\quad[A_{1},A_{3}]=A_{4},\quad[A_{1},A_{4}]=-A_{3},\quad[A_{1},A_{5}]=0,$
$[A_{2},A_{3}]=A_{5},\quad[A_{2},A_{4}]=0,\quad[A_{2},A_{5}]=-A_{3},$
$[A_{3},A_{4}]=A_{5},\quad[A_{3},A_{5}]=-A_{4},\quad[A_{4},A_{5}]=A_{3}.$
Second, all algebras possess nontrivial centers $Z(\mathcal{A})$, defined as the set of elements that commute with all others. The dimensionality of each center equals the number of $\hat{G}_{i_{\sigma}j_{\tau}}^{a_{\mu}b_{\nu}}$ terms in the symmetrized generator (i.e., 2, 4, and 6). All central elements are of the form $P[\{\hat{n}(q_{\sigma})\}]\hat{G}_{i_{\sigma}j_{\tau}}^{a_{\mu}b_{\nu}}$, where $P[\{\hat{n}(q_{\sigma})\}]$ are projector polynomials of occupation numbers with eigenvalues $\{0,1\}$. In the seniority-2 case, there are two central elements, $Z_{1}=A_{1}-A_{5}$ and $Z_{2}=A_{2}+A_{4}$, such that $Z(\mathcal{A}_{S2})=\mathrm{span}\{Z_{1},Z_{2}\}\cong\mathbb{R}^{2}$. Their explicit forms are
$Z_{1}$ $=\big(\hat{n}(j_{\beta})-\hat{n}(a_{\alpha})\big)^{2}\,\hat{G}_{i_{\alpha}i_{\beta}}^{\,j_{\alpha}a_{\beta}},$
$Z_{2}$ $=\big(\hat{n}(j_{\alpha})-\hat{n}(a_{\beta})\big)^{2}\,\hat{G}_{i_{\alpha}i_{\beta}}^{\,j_{\beta}a_{\alpha}}.$
Finally, each Lie algebra can be partitioned as a direct sum of its center and a semisimple part, which itself decomposes into simple ideals. For the seniority-2 case, the subspace $\mathfrak{s}:=\mathrm{span}\{A_{3},A_{4},A_{5}\}$ is an ideal and is isomorphic to $\mathfrak{so}(3)\cong\mathfrak{s}.$ Thus, the full Lie algebra can be expressed as $\mathcal{A}_{S2}=\mathfrak{s}\oplus Z(\mathcal{A}_{S2})\cong\mathfrak{so}(3)\oplus\mathbb{R}^{2}.$ Although this partitioning does not reduce the number of terms in Eq. (1), it facilitates the construction of the decomposition in the adjoint representation.
II.3 Adjoint representation
To determine the parameters in Eq. (1) we use the adjoint representation, which maps each algebra element to a matrix of dimension equal to that of the Lie algebra. For the algebras considered here, the dimensions (depending on seniority and index pattern) are $5$, $28$, and $84$. These sizes are substantially smaller than those arising in faithful representations built in the computational basis of the corresponding spin orbitals. For example, using the computational basis in the Fock space of the involved spin orbitals requires matrices of dimensions $64$ and $256$ for seniorities $2$ and $4$, respectively. Therefore, we adopt the adjoint representation, in which operator matrices are constructed from the commutator (structure) constants: each element $A_{i}$ corresponds to a matrix $M^{(i)}$ defined by
$[A_{i},A_{j}]=\sum_{k}M_{jk}^{(i)}A_{k},$ (12)
so that the entries of $M^{(i)}$ are the structure constants of the Lie algebra.
This representation is faithful on the Lie algebra modulo its center: all central elements map to the zero matrix because they commute with every element. Since all algebras studied here have nontrivial centers, we separate the central and semisimple contributions in Eq. (1) before applying the adjoint representation. The additional commutativity within the center provides a universal reduction that isolates the semisimple component, allowing us to carry out the decomposition entirely in the adjoint representation of the semisimple part.
For all three algebras, the initial unitary can be written as
$U(\theta)=\exp\left[\theta\sum_{i=1}^{K}c_{i}A_{i}\right],$ (13)
where $c_{i}$ are fixed coefficients [see Eqs. (3), (5), and (II.2)], and $K$ is the number of central elements in the algebra. Each central element can be written as $Z_{i}=A_{i}+S_{i}$, where $S_{i}=\sum_{s}d_{is}A_{s}$ involves only elements from the semisimple part. In general, the semisimple elements do not mutually commute, but all elements forming the $Z_{i}$ commute with one another. Using this commutativity, we obtain
$\exp\left[\theta\sum_{i=1}^{K}c_{i}A_{i}\right]$ $=$ $\exp\left[\theta\sum_{i=1}^{K}c_{i}(Z_{i}-S_{i})\right]$ (14)
$=$ $\exp\left[\theta\sum_{i=1}^{K}c_{i}Z_{i}\right]\exp\left[-\theta\sum_{i=1}^{K}c_{i}S_{i}\right].$ (15)
The central part then factors as
$\exp\left[\theta\sum_{i=1}^{K}c_{i}Z_{i}\right]=\prod_{i}e^{\theta c_{i}Z_{i}},$ (16)
and, since each $Z_{i}$ reduces to a linear combination of mutually commuting Pauli products, these exponentials are straightforward to implement. The combined semisimple contribution involves only elements $\{A_{s}\}$ from the semisimple part,
$U_{S}(\theta)=\exp\left[-\theta\sum_{i=1}^{K}c_{i}S_{i}\right]=\prod_{s}e^{t_{s}(\theta)A_{s}}=U_{S2}(\mathbf{t}).$ (17)
Because the adjoint representation is faithful on the semisimple subalgebra, we use it to fit the parameters $t_{s}(\theta)$ numerically.
We determine $t_{s}(\theta)$ by minimizing the Frobenius norm of the difference between the two sides of Eq. (17), evaluated via matrix exponentials:
$\mathbf{t}^{\star}(\theta)=\arg\min_{\mathbf{t}\in\mathbb{R}^{m}}\;\big\|U_{S}(\theta)-U_{S2}(\mathbf{t})\big\|_{F}^{2}.$ (18)
This objective is representation-invariant within $\exp(\mathfrak{s})$ and encodes the Baker–Campbell–Hausdorff identity for the chosen semisimple Lie algebra. In practice, we employ a smooth quasi-Newton optimizer (BFGS). Because $m\leq 24$ in all cases considered, each evaluation of the norm difference is inexpensive and convergence is rapid. Code to perform this minimisation is available on GitHub[8].
II.4 Implementation of unitaries
All semisimple generators ($A_{s}$) and central generators ($Z_{i}$) in Eqs. (17) and (16) map, under a fermion-to-qubit transform, to linear combinations of commuting Pauli products within each factor. For definiteness, we use the Jordan–Wigner (JW) mapping. Given $\mathbf{t}(\theta)$, each factor $\exp\!\big(t_{s}(\theta)A_{s}\big)$ in Eq. (17) and each central factor $\exp(\theta c_{i}Z_{i})$ in Eq. (16) can be expressed as a product of exponentials of individual Pauli products. Each commuting set is implemented by a standard sequence: basis change, CNOT ladder, $R_{z}$ rotation, and inverse ladder. Table 1 reports Pauli-string counts under JW for the factorizations described.
$\hat{G}^{\text{SA}}_{iiab}$ $\hat{G}^{\text{SA}}_{ijaa}$ $\hat{G}^{\text{SA}}_{ijab}$ (S) $\hat{G}^{\text{SA}^{\prime}}_{ijab}$ (T)
$N_{\text{Pauli}}^{\text{JW}}$ 48 48 384 640
Table 1: Number of Pauli strings in the factorized implementation under the JW mapping.
To reduce the Pauli-string count, the generators ($A_{s}$) are grouped into groups of commuting operators. For all of the three cases, seniority-2, seniority-4 (intermediate singlet), and, seniority-4 (intermediate triplet), this results in three groups. The groups of commuting generators for the seniority-2 case are,
$s$ $=\{1,5\},\quad s=\{2,4\},\quad s=\{3\}$ (19)
The groups for the seniority-4 cases can be found in Appendices AB.
III Results
III.1 Computational details
To evaluate the effect of spin-adapted fermionic generators on the efficiency and accuracy of variational algorithms, we carried out a series of adaptive VQE calculations using different operator pools. All calculations employed the STO-3G[7, 6] basis set and were converged when the maximum absolute value of the gradient of the operator pool was below $10^{-5}$ a.u. The state-vector simulator implemented in SlowQuant[9] was used to perform noiseless classical emulations. Energy minimizations employed the BFGS algorithm as implemented in SciPy[16].
The operator pool for the conventional fermionic singles and doubles (SD) adaptive fermionic ansatz consisted of $\{\hat{G}_{i_{\sigma}}^{a_{\sigma}},\,\hat{G}_{i_{\sigma}j_{\tau}}^{a_{\sigma}b_{\tau}},\,\hat{G}_{i_{\sigma}j_{\tau}}^{a_{\tau}b_{\sigma}}\},\quad i,j<a,b$, this ansatz is equal to the ADAPT ansatz[5]. For the spin-adapted fermionic ansatz ($\mathrm{S}_{\mathrm{SA}}\mathrm{D}_{\mathrm{SA}}$), the operator pool was $\{\hat{G}^{\mathrm{SA}}_{ia},\,\hat{G}^{\mathrm{SA}}_{ijab},\,\hat{G}^{\mathrm{SA}^{\prime}}_{ijab}\},\quad i,j<a,b$. Here, $\hat{G}^{\mathrm{SA}}_{ia}=\hat{G}_{i_{\alpha}}^{a_{\alpha}}+\hat{G}_{i_{\beta}}^{a_{\beta}}$, is a fermionic spin-adapted single excitation operator. Finally, for the spin-adapted singles and pair-doubles ansatz ($\mathrm{S}_{\mathrm{SA}}\mathrm{D}_{\mathrm{pair}}$), the operator pool was $\{\hat{G}^{\mathrm{SA}}_{ia},\,\hat{G}^{\mathrm{SA}}_{iiaa}\},\quad i<a,$ where $\hat{G}^{\mathrm{SA}}_{iiaa}$ corresponds to $\hat{G}_{i_{\alpha}i_{\beta}}^{a_{\alpha}a_{\beta}}$. Each adaptive VQE ansatz was constructed iteratively by selecting the operator with the largest energy gradient magnitude at each step until the convergence criterion was reached.
III.2 Adaptive VQE calculations
With these operator pools defined, we now compare their performance for representative molecular systems. Using the fermionic operator pool $\{\hat{G}_{i_{\sigma}}^{a_{\sigma}},\hat{G}_{i_{\sigma}j_{\tau}}^{a_{\sigma}b_{\tau}},\hat{G}_{i_{\sigma}j_{\tau}}^{a_{\tau}b_{\sigma}}\}$ constrains the variational space to determinants that conserve particle number and spin projection. In contrast, the spin-adapted fermionic operator pool $\{\hat{G}^{\mathrm{SA}}_{ia},\hat{G}^{\mathrm{SA}}_{ijab},\hat{G}^{\mathrm{SA}^{\prime}}_{ijab}\}$ further enforces total spin conservation, effectively spanning the configuration-state function (CSF) space. Because the nonrelativistic, spin-free electronic Hamiltonian commutes with the total spin operator, it is expected that using spin-adapted generators should reduce the number of variational parameters required for convergence.
To quantify the effect of spin adaptation on circuit compactness, we compared the number of variational parameters required for convergence in adaptive VQE expansions for two systems: H2O and BeH2, both in the STO-3G basis set.
Figure 1: Adaptive VQE calculations for $\text{H}_{2}\text{O}$ with $\theta_{\mathrm{HOH}}=109^{\circ}$ and $R_{\mathrm{OH}}=0.977$ Å. SD denotes the fermionic operator pool, $\text{S}_{\text{SA}}\text{D}_{\text{SA}}$ the spin-adapted fermionic operator pool, and $\text{S}_{\text{SA}}\text{D}_{\text{pair}}$ the pool of spin-adapted singles and pair doubles.
Figure 1 shows the convergence of adaptive VQE calculations for H2O relative to the full configuration interaction (FCI) energy. Using the fermionic operator pool (blue dashed line) required 104 variational parameters to reach the gradient threshold of $10^{-5}$ a.u. In contrast, the spin-adapted fermionic operator pool (red dashed line) achieved convergence with only 54 parameters, demonstrating that enforcing total spin conservation leads to a more compact ansatz. Physically, this reduction arises because spin-adapted generators only couple configurations within a single spin manifold, eliminating redundant directions in parameter space.
The alternative pool of spin-adapted singles and pair doubles (green dashed line) also conserves total spin and consists of circuits that can be directly implemented with known gate structures. However, this ansatz frequently becomes trapped in local minima and, in this case, converged to a solution $8.8\times 10^{-3}$ Hartree above the FCI reference. A similar shortcoming of pair-doubles-based operator pools has been reported previously by Burton et al.[3]
Figure 2: Adaptive VQE calculations for $\text{Be}\text{H}_{2}$ with $\theta_{\mathrm{HBeH}}=180^{\circ}$ and $R_{\mathrm{BeH}}=1.35$ Å. SD denotes the fermionic operator pool, $\text{S}_{\text{SA}}\text{D}_{\text{SA}}$ the spin-adapted fermionic operator pool, and $\text{S}_{\text{SA}}\text{D}_{\text{pair}}$ the pool of spin-adapted singles and pair doubles.
Figure 2 presents analogous results for BeH2. The same qualitative trend is observed: the fermionic operator pool required 95 parameters to converge, whereas the spin-adapted fermionic pool reached convergence with only 55 parameters. Again, the $\text{S}_{\text{SA}}\text{D}_{\text{pair}}$ operator pool became trapped in a local minimum and failed to reach the FCI energy. These results confirm that incorporating total spin symmetry systematically reduces the variational dimension required for accurate convergence.
An important advantage of using spin-adapted operators is their ability to prevent variational collapse to states of different total spin. To illustrate this point, we investigated the O2 molecule in its triplet configuration, where the risk of spin contamination is high.
The test system used the molecular orbitals obtained from a restricted open-shell Hartree–Fock (ROHF) triplet calculation. The reference state was chosen as the triplet CSF
$\frac{1}{\sqrt{2}}\big(\pi^{*}_{x}(\downarrow)\pi^{*}_{y}(\uparrow)+\pi^{*}_{x}(\uparrow)\pi^{*}_{y}(\downarrow)\big),$
corresponding to $m_{s}=0$.
Figure 3: Adaptive VQE calculations for $\text{O}_{2}$ triplet. SD denotes the fermionic operator pool, $\text{S}_{\text{SA}}\text{D}_{\text{SA}}$ the spin-adapted fermionic operator pool. FCI triplet and singlet reference curves are shown for comparison.
Figure 3 displays the potential energy curve of O2 for bond lengths between 1.8 and 2.4 Å. When using the spin-adapted fermionic operator pool ($\text{S}_{\text{SA}}\text{D}_{\text{SA}}$, red circles), the adaptive VQE energies follow the FCI triplet curve (black line) across all bond lengths. In contrast, using the conventional fermionic operator pool (SD, blue diamonds) leads to a collapse to the lower-energy singlet solution (green curve), even when starting from a triplet reference. This behavior arises because the SD operators do not conserve total spin and thus allow transitions between spin manifolds.
Previous studies have attempted to address this problem by adding penalty terms to the Hamiltonian during variational optimization[10] or by applying spin-projection techniques[15]. However, these approaches increase computational cost either by complicating the energy functional or by requiring additional ancilla qubits. Using spin-adapted operators naturally preserves the desired spin symmetry without modifying the Hamiltonian or introducing extra qubits, offering a simple and physically consistent solution.
IV Conclusions
We have introduced a framework for factorizing spin-adapted unitary transformations into ordered products of exponentials within compact Lie algebras of symmetrized excitations. By employing the adjoint representation, the exponential of a sum of spin-adapted generators can be obtained via a low-dimensional nonlinear optimization over matrix exponentials, yielding numerically exact amplitudes without recourse to symbolic Baker–Campbell–Hausdorff expansions. This formulation preserves total spin and particle-number symmetries by construction and produces efficient, low-depth circuit realizations suitable for use in variational adaptive algorithms such as adaptive VQE schemes.
The accompanying analysis of seniority-specific Lie algebras clarifies why such compact factorizations exist and how their algebraic structure influences parameter organization and numerical stability. In particular, the identification of small, closed algebras within each seniority sector accelerates convergence in the adjoint-space optimization and provides a foundation for constructing numerically stable, symmetry-consistent amplitude parametrizations. In benchmark adaptive VQE calculations, enforcing spin adaptation was shown to reduce the number of variational parameters by roughly a factor of two while maintaining exact spin symmetry throughout the optimization.
Building on these results, the combined algebraic and numerical framework presented here establishes a general pathway for translating symmetry-adapted unitary transformations into implementable quantum circuits. Extending this approach to higher-seniority, spin–orbit-coupled, or alternative symmetry-adapted spaces may enable the design of new ansätze with greater expressivity at reduced circuit depth. More broadly, the adjoint-based formulation offers a principled route linking Lie-algebraic structure to variational algorithm design, providing a systematic strategy for constructing efficient, symmetry-preserving transformations for quantum simulation.
Acknowledgments
A.F.I. acknowledge financial support from the Natural Sciences and Engineering Council of Canada (NSERC) and the Mønsted Foundation for his research stay at the Department of Chemistry, Technical University of Denmark (DTU). E.R.K. acknowledge the support of the Novo Nordisk Foundation (NNF) for the focused research project “Hybrid Quantum Chemistry on Hybrid Quantum Computers” (grant number NNFSA220080996).
Appendix A Seniority 4, intermediate singlet
This algebra is generated by $A_{1}=\hat{G}^{a_{\alpha}b_{\beta}}_{i_{\alpha}j_{\beta}}$, $A_{2}=\hat{G}^{a_{\beta}b_{\alpha}}_{i_{\beta}j_{\alpha}}$, $A_{3}=\hat{G}^{a_{\alpha}b_{\beta}}_{i_{\beta}j_{\alpha}}$, $A_{4}=\hat{G}^{a_{\beta}b_{\alpha}}_{i_{\alpha}j_{\beta}}$ from the symmetrized generator of Eq. (5). The generated Lie algebra has dimension $28$, $\mathcal{A}_{S4S}=\{A_{p}\}_{p=1}^{28}$, with the following generators:
First-Level Commutators:
$A_{5}$ $=\left(1-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})\right)\hat{G}_{i_{\alpha}j_{\beta}}^{i_{\beta}j_{\alpha}}$
$A_{6}$ $=\left(1-\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})\right)\hat{G}_{a_{\alpha}b_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{7}$ $=-\left(1-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})\right)\hat{G}_{a_{\alpha}b_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{8}$ $=-\left(1-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})\right)\hat{G}_{i_{\alpha}j_{\beta}}^{i_{\beta}j_{\alpha}}$
Second-Level Commutators:
$A_{9}$ $=-\left(1-\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})+2\hat{n}(i_{\alpha})\hat{n}(j_{\beta})\right)\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{10}$ $=-\left(1-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})+2\hat{n}(a_{\alpha})\hat{n}(b_{\beta})\right)\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{11}$ $=\left(1-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(i_{\beta})\hat{n}(b_{\beta})\right$.
$\quad\left.+\hat{n}(i_{\beta})\hat{n}(a_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(b_{\beta})+\hat{n}(j_{\alpha})\hat{n}(a_{\alpha})\right)\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{12}$ $=\left(1-\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})+\hat{n}(i_{\alpha})\hat{n}(a_{\beta})\right$.
$\quad\left.+\hat{n}(i_{\alpha})\hat{n}(b_{\alpha})+\hat{n}(j_{\beta})\hat{n}(a_{\beta})+\hat{n}(j_{\beta})\hat{n}(b_{\alpha})\right)\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{13}$ $=-\left(1-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})+2\hat{n}(a_{\beta})\hat{n}(b_{\alpha})\right)\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{14}$ $=-\left(1-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})+2\hat{n}(i_{\beta})\hat{n}(j_{\alpha})\right)\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{15}$ $=\left(1-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})+2\hat{n}(i_{\beta})\hat{n}(j_{\alpha})\right)\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{16}$ $=-\left(1-\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(i_{\alpha})\hat{n}(a_{\alpha})\right$.
$\quad\left.+\hat{n}(i_{\alpha})\hat{n}(b_{\beta})+\hat{n}(j_{\beta})\hat{n}(a_{\alpha})+\hat{n}(j_{\beta})\hat{n}(b_{\beta})\right)\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{17}$ $=\left(1-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})+2\hat{n}(a_{\alpha})\hat{n}(b_{\beta})\right)\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{18}$ $=-\left(1-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})+\hat{n}(i_{\beta})\hat{n}(a_{\beta})\right$.
$\quad\left.+\hat{n}(i_{\beta})\hat{n}(b_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(a_{\beta})+\hat{n}(j_{\alpha})\hat{n}(b_{\alpha})\right)\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{19}$ $=\left(1-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})+2\hat{n}(a_{\beta})\hat{n}(b_{\alpha})\right)\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{20}$ $=\left(1-\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})+2\hat{n}(i_{\alpha})\hat{n}(j_{\beta})\right)\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
Third-Level Commutators:
$A_{21}$ $=\left(1-\hat{n}(i_{\alpha})-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(j_{\beta})+\hat{n}(i_{\alpha})\hat{n}(i_{\beta})+\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})\right$.
$\quad\left.+\hat{n}(i_{\beta})\hat{n}(j_{\beta})+\hat{n}(j_{\alpha})\hat{n}(j_{\beta})+2\hat{n}(i_{\alpha})\hat{n}(j_{\beta})-2\hat{n}(i_{\alpha})\hat{n}(i_{\beta})\hat{n}(j_{\beta})\right$.
$\quad\left.-2\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})\hat{n}(j_{\beta})\right)\hat{G}_{a_{\alpha}b_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{22}$ $=\left(1-\hat{n}(a_{\alpha})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(a_{\alpha})\hat{n}(a_{\beta})+\hat{n}(a_{\alpha})\hat{n}(b_{\alpha})\right$.
$\quad\left.+\hat{n}(a_{\beta})\hat{n}(b_{\beta})+\hat{n}(b_{\alpha})\hat{n}(b_{\beta})+2\hat{n}(a_{\alpha})\hat{n}(b_{\beta})-2\hat{n}(a_{\alpha})\hat{n}(a_{\beta})\hat{n}(b_{\beta})\right$.
$\quad\left.-2\hat{n}(a_{\alpha})\hat{n}(b_{\alpha})\hat{n}(b_{\beta})\right)\hat{G}_{i_{\alpha}j_{\beta}}^{i_{\beta}j_{\alpha}}$
$A_{23}$ $=-\left(1-\hat{n}(a_{\alpha})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(a_{\alpha})\hat{n}(a_{\beta})+\hat{n}(a_{\alpha})\hat{n}(b_{\alpha})\right$.
$\quad\left.+\hat{n}(a_{\beta})\hat{n}(b_{\beta})+\hat{n}(b_{\alpha})\hat{n}(b_{\beta})+2\hat{n}(a_{\beta})\hat{n}(b_{\alpha})-2\hat{n}(a_{\alpha})\hat{n}(a_{\beta})\hat{n}(b_{\alpha})\right$.
$\quad\left.-2\hat{n}(a_{\beta})\hat{n}(b_{\alpha})\hat{n}(b_{\beta})\right)\hat{G}_{i_{\alpha}j_{\beta}}^{i_{\beta}j_{\alpha}}$
$A_{24}$ $=-\left(1-\hat{n}(i_{\alpha})-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(j_{\beta})+\hat{n}(i_{\alpha})\hat{n}(i_{\beta})+\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})\right$.
$\quad\left.+\hat{n}(i_{\beta})\hat{n}(j_{\beta})+\hat{n}(j_{\alpha})\hat{n}(j_{\beta})+2\hat{n}(i_{\beta})\hat{n}(j_{\alpha})-2\hat{n}(i_{\alpha})\hat{n}(i_{\beta})\hat{n}(j_{\alpha})\right$.
$\quad\left.-2\hat{n}(i_{\beta})\hat{n}(j_{\alpha})\hat{n}(j_{\beta})\right)\hat{G}_{a_{\alpha}b_{\beta}}^{a_{\beta}b_{\alpha}}$
Fourth-Level Commutators:
$A_{25}$ $=\left(1-\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})+\hat{n}(i_{\alpha})\hat{n}(b_{\alpha})+\hat{n}(i_{\alpha})\hat{n}(a_{\beta})\right$.
$\quad\left.+\hat{n}(j_{\beta})\hat{n}(b_{\alpha})+\hat{n}(j_{\beta})\hat{n}(a_{\beta})+2\hat{n}(i_{\alpha})\hat{n}(j_{\beta})+2\hat{n}(a_{\beta})\hat{n}(b_{\alpha})\right$.
$\quad\left.-2\hat{n}(i_{\alpha})\hat{n}(a_{\beta})\hat{n}(b_{\alpha})-2\hat{n}(j_{\beta})\hat{n}(a_{\beta})\hat{n}(b_{\alpha})-2\hat{n}(i_{\alpha})\hat{n}(j_{\beta})\hat{n}(b_{\alpha})\right$.
$\quad\left.-2\hat{n}(i_{\alpha})\hat{n}(j_{\beta})\hat{n}(a_{\beta})+4\hat{n}(i_{\alpha})\hat{n}(j_{\beta})\hat{n}(a_{\beta})\hat{n}(b_{\alpha})\right)\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{26}$ $=\left(1-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(j_{\alpha})\hat{n}(a_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(b_{\beta})\right$.
$\quad\left.+\hat{n}(i_{\beta})\hat{n}(a_{\alpha})+\hat{n}(i_{\beta})\hat{n}(b_{\beta})+2\hat{n}(i_{\beta})\hat{n}(j_{\alpha})+2\hat{n}(a_{\alpha})\hat{n}(b_{\beta})\right$.
$\quad\left.-2\hat{n}(i_{\beta})\hat{n}(j_{\alpha})\hat{n}(a_{\alpha})-2\hat{n}(i_{\beta})\hat{n}(j_{\alpha})\hat{n}(b_{\beta})-2\hat{n}(j_{\alpha})\hat{n}(a_{\alpha})\hat{n}(b_{\beta})\right$.
$\quad\left.-2\hat{n}(i_{\beta})\hat{n}(a_{\alpha})\hat{n}(b_{\beta})+4\hat{n}(i_{\beta})\hat{n}(j_{\alpha})\hat{n}(a_{\alpha})\hat{n}(b_{\beta})\right)\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{27}$ $=-\left(1-\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(i_{\alpha})\hat{n}(a_{\alpha})+\hat{n}(i_{\alpha})\hat{n}(b_{\beta})\right$.
$\quad\left.+\hat{n}(j_{\beta})\hat{n}(a_{\alpha})+\hat{n}(j_{\beta})\hat{n}(b_{\beta})+2\hat{n}(i_{\alpha})\hat{n}(j_{\beta})+2\hat{n}(a_{\alpha})\hat{n}(b_{\beta})\right$.
$\quad\left.-2\hat{n}(i_{\alpha})\hat{n}(a_{\alpha})\hat{n}(b_{\beta})-2\hat{n}(j_{\beta})\hat{n}(a_{\alpha})\hat{n}(b_{\beta})-2\hat{n}(i_{\alpha})\hat{n}(j_{\beta})\hat{n}(a_{\alpha})\right$.
$\quad\left.-2\hat{n}(i_{\alpha})\hat{n}(j_{\beta})\hat{n}(b_{\beta})+4\hat{n}(i_{\alpha})\hat{n}(j_{\beta})\hat{n}(a_{\alpha})\hat{n}(b_{\beta})\right)\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{28}$ $=-\left(1-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})+\hat{n}(i_{\beta})\hat{n}(a_{\beta})+\hat{n}(i_{\beta})\hat{n}(b_{\alpha})\right$.
$\quad\left.+\hat{n}(j_{\alpha})\hat{n}(a_{\beta})+\hat{n}(j_{\alpha})\hat{n}(b_{\alpha})+2\hat{n}(i_{\beta})\hat{n}(j_{\alpha})+2\hat{n}(a_{\beta})\hat{n}(b_{\alpha})\right$.
$\quad\left.-2\hat{n}(i_{\beta})\hat{n}(a_{\beta})\hat{n}(b_{\alpha})-2\hat{n}(j_{\alpha})\hat{n}(a_{\beta})\hat{n}(b_{\alpha})-2\hat{n}(i_{\beta})\hat{n}(j_{\alpha})\hat{n}(a_{\beta})\right$.
$\quad\left.-2\hat{n}(i_{\beta})\hat{n}(j_{\alpha})\hat{n}(b_{\alpha})+4\hat{n}(i_{\beta})\hat{n}(j_{\alpha})\hat{n}(a_{\beta})\hat{n}(b_{\alpha})\right)\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
The commutation relations of these are given in Table 2.
Commuting groups
The groups of commuting generators are,
$s$ $=\{1,2,15,16,17,18,19,20,27,28\}$ (20)
$s$ $=\{3,4,9,10,11,12,13,14,25,26\}$ (21)
$s$ $=\{5,6,7,8,21,22,23,24\}$ (22)
Table 2: Commutation Table for the 28-Element Seniority-4 Lie Algebra $A_{i},A_{j}]$
[i, j] 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28
1 0 0 5 6 9 10 11 12 -5 -6 21 22 23 24 0 0 0 0 0 0 -11 -12 25 26 -23 -24 0 0
2 0 0 7 8 11 12 13 14 21 22 23 24 -7 -8 0 0 0 0 0 0 25 26 -11 -12 -21 -22 0 0
3 -5 -7 0 0 15 16 17 18 0 0 0 0 0 0 -5 24 -7 22 23 21 27 -18 28 -16 0 0 -21 -23
4 -6 -8 0 0 16 19 18 20 0 0 0 0 0 0 24 23 22 21 -6 -8 -18 27 -16 28 0 0 -22 -24
5 -9 -11 -15 -16 0 0 0 0 15 16 -27 18 -28 16 -9 26 -11 -12 25 -11 0 0 0 0 28 -16 11 -25
6 -10 -12 -16 -19 0 0 0 0 16 19 18 -27 16 -28 26 25 -12 -11 -10 -12 0 0 0 0 -16 28 12 -26
7 -11 -13 -17 -18 0 0 0 0 -27 18 -28 16 17 18 -11 -12 -13 26 -11 25 0 0 0 0 27 -18 -25 11
8 -12 -14 -18 -20 0 0 0 0 18 -27 16 -28 18 20 -12 -11 26 25 -12 -14 0 0 0 0 -18 27 -26 12
9 5 -21 0 0 -15 -16 27 -18 0 0 0 0 0 0 5 -24 -21 -22 -23 -21 -27 18 -28 16 0 0 21 23
10 6 -22 0 0 -16 -19 -18 27 0 0 0 0 0 0 -24 -23 -22 -21 6 -22 18 -27 16 -28 0 0 22 24
11 -21 -23 0 0 27 -18 28 -16 0 0 0 0 0 0 -21 -22 -23 -24 -21 -23 -28 16 -27 18 0 0 23 21
12 -22 -24 0 0 -18 27 -16 28 0 0 0 0 0 0 -22 -21 -24 -23 -22 -24 16 -28 18 -27 0 0 24 22
13 -23 7 0 0 28 -16 -17 -18 0 0 0 0 0 0 -23 -24 7 -22 -23 -21 -27 18 -28 16 0 0 21 23
14 -24 8 0 0 -16 28 -18 -20 0 0 0 0 0 0 -24 -23 -22 -21 -24 8 18 -27 16 -28 0 0 22 24
15 0 0 5 -24 9 -26 11 12 -5 24 21 22 23 24 0 0 0 0 0 0 -11 -12 25 26 -23 -24 0 0
16 0 0 -24 -23 -26 -25 12 11 24 23 22 21 24 23 0 0 0 0 0 0 -12 -11 26 25 -24 -23 0 0
17 0 0 7 -22 11 12 13 -26 21 22 23 24 -7 22 0 0 0 0 0 0 25 26 -11 -12 -21 -22 0 0
18 0 0 -22 -21 12 11 -26 -25 22 21 24 23 22 21 0 0 0 0 0 0 26 25 -12 -11 -22 -21 0 0
19 0 0 -23 6 -25 10 11 12 23 -6 21 22 23 24 0 0 0 0 0 0 -11 -12 25 26 -23 -24 0 0
20 0 0 -21 8 11 12 -25 14 21 22 23 24 21 -8 0 0 0 0 0 0 25 26 -11 -12 -21 -22 0 0
21 11 -25 -27 18 0 0 0 0 27 -18 28 -16 27 -18 11 12 -25 -26 11 -25 0 0 0 0 -27 18 25 -11
22 12 -26 18 -27 0 0 0 0 -18 27 -16 28 -18 27 12 11 -26 -25 12 -26 0 0 0 0 18 -27 26 -12
23 -25 11 -28 16 0 0 0 0 28 -16 27 -18 28 -16 -25 -26 11 12 -25 11 0 0 0 0 -28 16 -11 25
24 -26 12 16 -28 0 0 0 0 -16 28 -18 27 -16 28 -26 -25 12 11 -26 12 0 0 0 0 16 -28 -12 26
25 23 21 0 0 -28 16 -27 18 0 0 0 0 0 0 23 24 21 22 23 21 27 -18 28 -16 0 0 -21 -23
26 24 22 0 0 16 -28 18 -27 0 0 0 0 0 0 24 23 22 21 24 22 -18 27 -16 28 0 0 -22 -24
27 0 0 21 22 -11 -12 25 26 -21 -22 -23 -24 -21 -22 0 0 0 0 0 0 -25 -26 11 12 21 22 0 0
28 0 0 23 24 25 26 -11 -12 -23 -24 -21 -22 -23 -24 0 0 0 0 0 0 11 12 -25 -26 23 24 0 0
Algebraic structure
Let us denote $\mathfrak{s}=[\mathcal{A}_{S4S},\mathcal{A}_{S4S}]$ the derived algebra. The restriction of the Killing form to $\mathfrak{s}$ is negative-definite and nondegenerate, i.e. $\mathfrak{s}$ is of compact type. Additionally, $\mathfrak{s}$ has no nontrivial proper ideals. From the commutator table, $\mathfrak{s}=\mathrm{span}\{A_{5},A_{6},\dots,A_{28}\}$, $\dim\mathfrak{s}=24$. Consequently, $\mathcal{A}_{S4S}$ admits a Levi decomposition $\mathcal{A}_{S4S}=\mathfrak{s}\oplus\mathfrak{r}$, $\mathfrak{r}=Z(\mathcal{A}_{S4S})$ solvable radical, $\dim=4$.
While no single $A_{i}$ is central, the center $Z(\mathcal{A}_{S4S})$ is $4$-dimensional, with the following convenient basis of linear combinations:
$Z_{1}$ $=A_{3}+A_{9}+A_{13}+A_{25},$
$Z_{2}$ $=A_{4}+A_{10}+A_{14}+A_{26},$
$Z_{3}$ $=-\,A_{2}+A_{17}+A_{20}+A_{27},$
$Z_{4}$ $=-\,A_{1}+A_{15}+A_{19}+A_{28},$
thus, $Z(\mathcal{A}_{S4S})=\mathrm{span}\{Z_{1},Z_{2},Z_{3},Z_{4}\}\cong\mathbb{R}^{4}$. The central directions have the following explicit forms:
$Z_{1}$ $=\big(\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})\big)^{2}\big(\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})\big)^{2}\,\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$Z_{2}$ $=\big(\hat{n}(j_{\alpha})-\hat{n}(i_{\beta})\big)^{2}\big(\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})\big)^{2}\,\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$Z_{3}$ $=-\Big(\hat{n}(i_{\alpha})\hat{n}(a_{\alpha})\big(1-\hat{n}(b_{\beta})\big)\big(1-\hat{n}(j_{\beta})\big)+\hat{n}(j_{\beta})\hat{n}(b_{\beta})\big(1-\hat{n}(i_{\alpha})\big)\big(1-\hat{n}(a_{\alpha})\big)\Big)\,\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$Z_{4}$ $=-\Big(\hat{n}(i_{\beta})\hat{n}(a_{\beta})\big(1-\hat{n}(b_{\alpha})\big)\big(1-\hat{n}(j_{\alpha})\big)+\hat{n}(j_{\alpha})\hat{n}(b_{\alpha})\big(1-\hat{n}(i_{\beta})\big)\big(1-\hat{n}(a_{\beta})\big)\Big)\,\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
For $i=1,2,3,4$ let $\mathcal{A}_{i}$ denote the $4$-tuple consisting of $A_{i}$ together with its three “partners”:
$\mathcal{A}_{1}$ $=\{A_{1},A_{15},A_{19},A_{28}\},$ $\mathcal{A}_{2}$ $=\{A_{2},A_{17},A_{20},A_{27}\},$
$\mathcal{A}_{3}$ $=\{A_{3},A_{9},A_{13},A_{25}\},$ $\mathcal{A}_{4}$ $=\{A_{4},A_{10},A_{14},A_{26}\}.$
Within each $\mathcal{A}_{i}$ all basis elements commute pairwise, so $\langle\mathcal{A}_{i}\rangle$ is a $4$-dimensional abelian subalgebra. However, only the specific linear combination displayed above lies in the center; the remaining vectors in each $\mathcal{A}_{i}$ are not central.
Appendix B Seniority 4, intermediate triplet
This 84-dimensional Lie algebra originates from the commutator closure of 6 elementary generators (the closure produces 84 linearly independent as fermionic operators)
$A_{1}$ $=\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{2}$ $=\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{3}$ $=\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{4}$ $=\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{5}$ $=\hat{G}_{i_{\alpha}j_{\alpha}}^{a_{\alpha}b_{\alpha}}$
$A_{6}$ $=\hat{G}_{i_{\beta}j_{\beta}}^{a_{\beta}b_{\beta}}$
In what follows we list all other generators of this algebra.
First-Level Commutators:
$A_{7}$ $=(1-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{i_{\beta}j_{\alpha}}$
$A_{8}$ $=(1-\hat{n}(i_{\alpha})-\hat{n}(j_{\beta}))\hat{G}_{a_{\alpha}b_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{9}$ $=(\hat{n}(i_{\alpha})-\hat{n}(a_{\alpha}))\hat{G}_{j_{\alpha}b_{\beta}}^{j_{\beta}b_{\alpha}}$
$A_{10}$ $=(\hat{n}(b_{\beta})-\hat{n}(j_{\beta}))\hat{G}_{i_{\alpha}a_{\beta}}^{i_{\beta}a_{\alpha}}$
$A_{11}$ $=-(1-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha}))\hat{G}_{a_{\alpha}b_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{12}$ $=-(1-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{i_{\beta}j_{\alpha}}$
$A_{13}$ $=(\hat{n}(j_{\alpha})-\hat{n}(b_{\alpha}))\hat{G}_{i_{\alpha}a_{\beta}}^{i_{\beta}a_{\alpha}}$
$A_{14}$ $=(\hat{n}(a_{\beta})-\hat{n}(i_{\beta}))\hat{G}_{j_{\alpha}b_{\beta}}^{j_{\beta}b_{\alpha}}$
$A_{15}$ $=(\hat{n}(j_{\alpha})-\hat{n}(a_{\alpha}))\hat{G}_{i_{\alpha}b_{\beta}}^{i_{\beta}b_{\alpha}}$
$A_{16}$ $=(\hat{n}(b_{\beta})-\hat{n}(i_{\beta}))\hat{G}_{j_{\alpha}a_{\beta}}^{j_{\beta}a_{\alpha}}$
$A_{17}$ $=(\hat{n}(i_{\alpha})-\hat{n}(b_{\alpha}))\hat{G}_{j_{\alpha}a_{\beta}}^{j_{\beta}a_{\alpha}}$
$A_{18}$ $=(\hat{n}(a_{\beta})-\hat{n}(j_{\beta}))\hat{G}_{i_{\alpha}b_{\beta}}^{i_{\beta}b_{\alpha}}$
Second-Level Commutators:
$A_{19}$ $=-(1-\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})+2\hat{n}(i_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{20}$ $=-(1-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})+2\hat{n}(a_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{21}$ $=-(\hat{n}(j_{\beta})+\hat{n}(b_{\beta})-2\hat{n}(j_{\beta})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\alpha}}^{a_{\alpha}b_{\alpha}}$
$A_{22}$ $=-(\hat{n}(i_{\alpha})+\hat{n}(a_{\alpha})-2\hat{n}(i_{\alpha})\hat{n}(a_{\alpha}))\hat{G}_{i_{\beta}j_{\beta}}^{a_{\beta}b_{\beta}}$
$A_{23}$ $=(1-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(i_{\beta})\hat{n}(b_{\beta})$
$\quad+\hat{n}(i_{\beta})\hat{n}(a_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(b_{\beta})+\hat{n}(j_{\alpha})\hat{n}(a_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{24}$ $=(1-\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})+\hat{n}(i_{\alpha})\hat{n}(a_{\beta})+\hat{n}(i_{\alpha})\hat{n}(b_{\alpha})$
$\quad+\hat{n}(j_{\beta})\hat{n}(a_{\beta})+\hat{n}(j_{\beta})\hat{n}(b_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{25}$ $=(\hat{n}(i_{\alpha})\hat{n}(b_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(a_{\alpha})-\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(b_{\alpha}))\hat{G}_{i_{\beta}j_{\beta}}^{a_{\beta}b_{\beta}}$
$A_{26}$ $=(\hat{n}(i_{\beta})\hat{n}(b_{\beta})+\hat{n}(j_{\beta})\hat{n}(a_{\beta})-\hat{n}(i_{\beta})\hat{n}(j_{\beta})-\hat{n}(a_{\beta})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\alpha}}^{a_{\alpha}b_{\alpha}}$
$A_{27}$ $=-(1-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})+2\hat{n}(a_{\beta})\hat{n}(b_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{28}$ $=-(1-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})+2\hat{n}(i_{\beta})\hat{n}(j_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{29}$ $=-(\hat{n}(i_{\beta})+\hat{n}(a_{\beta})-2\hat{n}(i_{\beta})\hat{n}(a_{\beta}))\hat{G}_{i_{\alpha}j_{\alpha}}^{a_{\alpha}b_{\alpha}}$
$A_{30}$ $=-(\hat{n}(j_{\alpha})+\hat{n}(b_{\alpha})-2\hat{n}(j_{\alpha})\hat{n}(b_{\alpha}))\hat{G}_{i_{\beta}j_{\beta}}^{a_{\beta}b_{\beta}}$
$A_{31}$ $=(1-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})+2\hat{n}(i_{\beta})\hat{n}(j_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{32}$ $=-(1-\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(i_{\alpha})\hat{n}(a_{\alpha})$
$\quad+\hat{n}(i_{\alpha})\hat{n}(b_{\beta})+\hat{n}(j_{\beta})\hat{n}(a_{\alpha})+\hat{n}(j_{\beta})\hat{n}(b_{\beta}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{33}$ $=(1-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})+2\hat{n}(a_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{34}$ $=-(1-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})+\hat{n}(i_{\beta})\hat{n}(a_{\beta})$
$\quad+\hat{n}(i_{\beta})\hat{n}(b_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(a_{\beta})+\hat{n}(j_{\alpha})\hat{n}(b_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{35}$ $=-(\hat{n}(i_{\beta})+\hat{n}(b_{\beta})-2\hat{n}(i_{\beta})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\alpha}}^{a_{\alpha}b_{\alpha}}$
$A_{36}$ $=-(\hat{n}(j_{\alpha})+\hat{n}(a_{\alpha})-2\hat{n}(j_{\alpha})\hat{n}(a_{\alpha}))\hat{G}_{i_{\beta}j_{\beta}}^{a_{\beta}b_{\beta}}$
$A_{37}$ $=(\hat{n}(i_{\alpha})\hat{n}(a_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(b_{\alpha})-\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(b_{\alpha}))\hat{G}_{i_{\beta}j_{\beta}}^{a_{\beta}b_{\beta}}$
$A_{38}$ $=(\hat{n}(i_{\beta})\hat{n}(a_{\beta})+\hat{n}(j_{\beta})\hat{n}(b_{\beta})-\hat{n}(i_{\beta})\hat{n}(j_{\beta})-\hat{n}(a_{\beta})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\alpha}}^{a_{\alpha}b_{\alpha}}$
$A_{39}$ $=(1-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})+2\hat{n}(a_{\beta})\hat{n}(b_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{40}$ $=(1-\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})+2\hat{n}(i_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{41}$ $=(\hat{n}(j_{\alpha})+\hat{n}(b_{\alpha})-2\hat{n}(j_{\alpha})\hat{n}(b_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{42}$ $=(\hat{n}(i_{\alpha})\hat{n}(j_{\beta})-\hat{n}(i_{\alpha})\hat{n}(b_{\beta})-\hat{n}(j_{\beta})\hat{n}(a_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{43}$ $=(\hat{n}(i_{\alpha})+\hat{n}(a_{\alpha})-2\hat{n}(i_{\alpha})\hat{n}(a_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{44}$ $=(\hat{n}(i_{\beta})\hat{n}(j_{\alpha})-\hat{n}(i_{\beta})\hat{n}(b_{\alpha})-\hat{n}(j_{\alpha})\hat{n}(a_{\beta})+\hat{n}(a_{\beta})\hat{n}(b_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{45}$ $=(\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})-2\hat{n}(i_{\alpha})\hat{n}(b_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{46}$ $=(\hat{n}(i_{\beta})\hat{n}(j_{\alpha})-\hat{n}(i_{\beta})\hat{n}(a_{\alpha})-\hat{n}(j_{\alpha})\hat{n}(b_{\beta})+\hat{n}(a_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{47}$ $=(\hat{n}(j_{\alpha})+\hat{n}(a_{\alpha})-2\hat{n}(j_{\alpha})\hat{n}(a_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{48}$ $=(\hat{n}(i_{\alpha})\hat{n}(j_{\beta})-\hat{n}(i_{\alpha})\hat{n}(a_{\beta})-\hat{n}(j_{\beta})\hat{n}(b_{\alpha})+\hat{n}(a_{\beta})\hat{n}(b_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
Third-Level Commutators:
$A_{49}$ $=(1-\hat{n}(i_{\alpha})-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(j_{\beta})+\hat{n}(i_{\alpha})\hat{n}(i_{\beta})+\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})+\hat{n}(i_{\beta})\hat{n}(j_{\beta})$
$\quad+\hat{n}(j_{\alpha})\hat{n}(j_{\beta})+2\hat{n}(i_{\alpha})\hat{n}(j_{\beta})-2\hat{n}(i_{\alpha})\hat{n}(i_{\beta})\hat{n}(j_{\beta})-2\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{a_{\alpha}b_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{50}$ $=(1-\hat{n}(a_{\alpha})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(a_{\alpha})\hat{n}(a_{\beta})+\hat{n}(a_{\alpha})\hat{n}(b_{\alpha})+\hat{n}(a_{\beta})\hat{n}(b_{\beta})$
$\quad+\hat{n}(b_{\alpha})\hat{n}(b_{\beta})+2\hat{n}(a_{\alpha})\hat{n}(b_{\beta})-2\hat{n}(a_{\alpha})\hat{n}(a_{\beta})\hat{n}(b_{\beta})-2\hat{n}(a_{\alpha})\hat{n}(b_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{i_{\beta}j_{\alpha}}$
$A_{51}$ $=(\hat{n}(j_{\beta})\hat{n}(b_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(b_{\beta})-\hat{n}(j_{\alpha})\hat{n}(j_{\beta})-\hat{n}(j_{\alpha})\hat{n}(b_{\beta})$
$\quad+2\hat{n}(j_{\alpha})\hat{n}(j_{\beta})\hat{n}(b_{\beta})-2\hat{n}(j_{\beta})\hat{n}(b_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}a_{\beta}}^{i_{\beta}a_{\alpha}}$
$A_{52}$ $=(\hat{n}(i_{\alpha})\hat{n}(i_{\beta})-\hat{n}(i_{\alpha})\hat{n}(a_{\beta})+\hat{n}(i_{\beta})\hat{n}(a_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(a_{\beta})$
$\quad-2\hat{n}(i_{\alpha})\hat{n}(i_{\beta})\hat{n}(a_{\alpha})+2\hat{n}(i_{\alpha})\hat{n}(a_{\alpha})\hat{n}(a_{\beta}))\hat{G}_{j_{\alpha}b_{\beta}}^{j_{\beta}b_{\alpha}}$
$A_{53}$ $=-(1-\hat{n}(a_{\alpha})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(a_{\alpha})\hat{n}(a_{\beta})+\hat{n}(a_{\alpha})\hat{n}(b_{\alpha})+\hat{n}(a_{\beta})\hat{n}(b_{\beta})$
$\quad+\hat{n}(b_{\alpha})\hat{n}(b_{\beta})+2\hat{n}(a_{\beta})\hat{n}(b_{\alpha})-2\hat{n}(a_{\alpha})\hat{n}(a_{\beta})\hat{n}(b_{\alpha})-2\hat{n}(a_{\beta})\hat{n}(b_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{i_{\beta}j_{\alpha}}$
$A_{54}$ $=-(1-\hat{n}(i_{\alpha})-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(j_{\beta})+\hat{n}(i_{\alpha})\hat{n}(i_{\beta})+\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})+\hat{n}(i_{\beta})\hat{n}(j_{\beta})$
$\quad+\hat{n}(j_{\alpha})\hat{n}(j_{\beta})+2\hat{n}(i_{\beta})\hat{n}(j_{\alpha})-2\hat{n}(i_{\alpha})\hat{n}(i_{\beta})\hat{n}(j_{\alpha})-2\hat{n}(i_{\beta})\hat{n}(j_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{a_{\alpha}b_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{55}$ $=(\hat{n}(a_{\alpha})\hat{n}(a_{\beta})-\hat{n}(i_{\alpha})\hat{n}(i_{\beta})-\hat{n}(i_{\alpha})\hat{n}(a_{\beta})+\hat{n}(i_{\beta})\hat{n}(a_{\alpha})$
$\quad+2\hat{n}(i_{\alpha})\hat{n}(i_{\beta})\hat{n}(a_{\beta})-2\hat{n}(i_{\beta})\hat{n}(a_{\alpha})\hat{n}(a_{\beta}))\hat{G}_{j_{\alpha}b_{\beta}}^{j_{\beta}b_{\alpha}}$
$A_{56}$ $=(\hat{n}(j_{\alpha})\hat{n}(j_{\beta})-\hat{n}(j_{\alpha})\hat{n}(b_{\beta})+\hat{n}(j_{\beta})\hat{n}(b_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(b_{\beta})$
$\quad-2\hat{n}(j_{\alpha})\hat{n}(j_{\beta})\hat{n}(b_{\alpha})+2\hat{n}(j_{\alpha})\hat{n}(b_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}a_{\beta}}^{i_{\beta}a_{\alpha}}$
$A_{57}$ $=-(\hat{n}(i_{\alpha})-\hat{n}(i_{\alpha})\hat{n}(i_{\beta})-\hat{n}(i_{\alpha})\hat{n}(a_{\alpha})-\hat{n}(i_{\beta})\hat{n}(a_{\alpha})+2\hat{n}(i_{\alpha})\hat{n}(i_{\beta})\hat{n}(a_{\alpha}))\hat{G}_{j_{\alpha}b_{\beta}}^{j_{\beta}b_{\alpha}}$
$A_{58}$ $=(\hat{n}(j_{\beta})-\hat{n}(j_{\alpha})\hat{n}(j_{\beta})-\hat{n}(j_{\alpha})\hat{n}(b_{\beta})-\hat{n}(j_{\beta})\hat{n}(b_{\beta})+2\hat{n}(j_{\alpha})\hat{n}(j_{\beta})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}a_{\beta}}^{i_{\beta}a_{\alpha}}$
$A_{59}$ $=(\hat{n}(b_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(b_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(b_{\beta})-\hat{n}(b_{\alpha})\hat{n}(b_{\beta})+2\hat{n}(a_{\alpha})\hat{n}(b_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{i_{\beta}j_{\alpha}}$
$A_{60}$ $=(\hat{n}(j_{\alpha})-\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(i_{\alpha})\hat{n}(j_{\beta})-\hat{n}(j_{\alpha})\hat{n}(j_{\beta})+2\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{a_{\alpha}b_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{61}$ $=(2\hat{n}(b_{\beta})\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})+\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})\hat{n}(j_{\alpha}))\hat{G}_{i_{\alpha}a_{\beta}}^{i_{\beta}a_{\alpha}}$
$A_{62}$ $=(\hat{n}(a_{\beta})\hat{n}(i_{\beta})-\hat{n}(a_{\beta})+\hat{n}(a_{\beta})\hat{n}(a_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(i_{\beta})-2\hat{n}(a_{\beta})\hat{n}(a_{\alpha})\hat{n}(i_{\beta}))\hat{G}_{j_{\alpha}b_{\beta}}^{j_{\beta}b_{\alpha}}$
$A_{63}$ $=(\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(i_{\beta})\hat{n}(i_{\alpha})-\hat{n}(i_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(i_{\beta})-2\hat{n}(j_{\alpha})\hat{n}(i_{\beta})\hat{n}(i_{\alpha}))\hat{G}_{a_{\alpha}b_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{64}$ $=(\hat{n}(b_{\alpha})\hat{n}(a_{\alpha})+\hat{n}(a_{\beta})\hat{n}(a_{\alpha})-\hat{n}(a_{\alpha})-2\hat{n}(b_{\alpha})\hat{n}(a_{\beta})\hat{n}(a_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(a_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{i_{\beta}j_{\alpha}}$
$A_{65}$ $=(\hat{n}(j_{\beta})\hat{n}(j_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(j_{\alpha})-2\hat{n}(a_{\alpha})\hat{n}(j_{\beta})\hat{n}(j_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{i_{\alpha}b_{\beta}}^{i_{\beta}b_{\alpha}}$
$A_{66}$ $=(\hat{n}(i_{\beta})-\hat{n}(i_{\beta})\hat{n}(i_{\alpha})-\hat{n}(b_{\beta})\hat{n}(i_{\beta})+2\hat{n}(b_{\beta})\hat{n}(i_{\beta})\hat{n}(i_{\alpha})-\hat{n}(b_{\beta})\hat{n}(i_{\alpha}))\hat{G}_{j_{\alpha}a_{\beta}}^{j_{\beta}a_{\alpha}}$
$A_{67}$ $=(\hat{n}(b_{\beta})\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(i_{\beta})\hat{n}(i_{\alpha})-\hat{n}(b_{\beta})\hat{n}(b_{\alpha})\hat{n}(i_{\beta})+\hat{n}(b_{\beta})\hat{n}(i_{\beta})\hat{n}(i_{\alpha})$
$\quad+\hat{n}(b_{\alpha})\hat{n}(i_{\beta})-\hat{n}(b_{\beta})\hat{n}(i_{\alpha}))\hat{G}_{j_{\alpha}a_{\beta}}^{j_{\beta}a_{\alpha}}$
$A_{68}$ $=-(\hat{n}(a_{\alpha})\hat{n}(j_{\beta})\hat{n}(j_{\alpha})-\hat{n}(a_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(a_{\beta})\hat{n}(j_{\beta})\hat{n}(j_{\alpha})+\hat{n}(a_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\beta})$
$\quad+\hat{n}(a_{\beta})\hat{n}(j_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{i_{\alpha}b_{\beta}}^{i_{\beta}b_{\alpha}}$
$A_{69}$ $=-(\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(a_{\alpha})-2\hat{n}(a_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})+\hat{n}(a_{\beta})\hat{n}(a_{\alpha})+\hat{n}(a_{\beta})\hat{n}(j_{\alpha}))\hat{G}_{i_{\alpha}b_{\beta}}^{i_{\beta}b_{\alpha}}$
$A_{70}$ $=(\hat{n}(b_{\beta})\hat{n}(i_{\beta})-\hat{n}(b_{\beta})-2\hat{n}(b_{\beta})\hat{n}(b_{\alpha})\hat{n}(i_{\beta})+\hat{n}(b_{\beta})\hat{n}(b_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(i_{\beta}))\hat{G}_{j_{\alpha}a_{\beta}}^{j_{\beta}a_{\alpha}}$
$A_{71}$ $=(2\hat{n}(b_{\beta})\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})\hat{n}(i_{\alpha}))\hat{G}_{j_{\alpha}a_{\beta}}^{j_{\beta}a_{\alpha}}$
$A_{72}$ $=(\hat{n}(a_{\beta})\hat{n}(j_{\beta})-\hat{n}(a_{\beta})+\hat{n}(a_{\beta})\hat{n}(a_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(j_{\beta})-2\hat{n}(a_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{i_{\alpha}b_{\beta}}^{i_{\beta}b_{\alpha}}$
$A_{73}$ $=(1-\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})+\hat{n}(i_{\alpha})\hat{n}(b_{\alpha})+\hat{n}(i_{\alpha})\hat{n}(a_{\beta})+\hat{n}(j_{\beta})\hat{n}(b_{\alpha})$
$\quad+\hat{n}(j_{\beta})\hat{n}(a_{\beta})+2\hat{n}(i_{\alpha})\hat{n}(j_{\beta})+2\hat{n}(a_{\beta})\hat{n}(b_{\alpha})-2\hat{n}(i_{\alpha})\hat{n}(a_{\beta})\hat{n}(b_{\alpha})-2\hat{n}(j_{\beta})\hat{n}(a_{\beta})\hat{n}(b_{\alpha})$
$\quad-2\hat{n}(i_{\alpha})\hat{n}(j_{\beta})\hat{n}(b_{\alpha})-2\hat{n}(i_{\alpha})\hat{n}(j_{\beta})\hat{n}(a_{\beta})+4\hat{n}(i_{\alpha})\hat{n}(j_{\beta})\hat{n}(a_{\beta})\hat{n}(b_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{74}$ $=(1-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(j_{\alpha})\hat{n}(a_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(b_{\beta})+\hat{n}(i_{\beta})\hat{n}(a_{\alpha})$
$\quad+\hat{n}(i_{\beta})\hat{n}(b_{\beta})+2\hat{n}(i_{\beta})\hat{n}(j_{\alpha})+2\hat{n}(a_{\alpha})\hat{n}(b_{\beta})-2\hat{n}(i_{\beta})\hat{n}(j_{\alpha})\hat{n}(a_{\alpha})-2\hat{n}(i_{\beta})\hat{n}(j_{\alpha})\hat{n}(b_{\beta})$
$\quad-2\hat{n}(j_{\alpha})\hat{n}(a_{\alpha})\hat{n}(b_{\beta})-2\hat{n}(i_{\beta})\hat{n}(a_{\alpha})\hat{n}(b_{\beta})+4\hat{n}(i_{\beta})\hat{n}(j_{\alpha})\hat{n}(a_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{75}$ $=(\hat{n}(a_{\beta})\hat{n}(j_{\beta})-2\hat{n}(b_{\beta})\hat{n}(j_{\beta})\hat{n}(i_{\beta})-2\hat{n}(b_{\beta})\hat{n}(a_{\beta})\hat{n}(j_{\beta})+\hat{n}(j_{\beta})\hat{n}(i_{\beta})$
$\quad+4\hat{n}(b_{\beta})\hat{n}(a_{\beta})\hat{n}(j_{\beta})\hat{n}(i_{\beta})-2\hat{n}(a_{\beta})\hat{n}(j_{\beta})\hat{n}(i_{\beta})+\hat{n}(b_{\beta})\hat{n}(i_{\beta})+\hat{n}(b_{\beta})\hat{n}(a_{\beta})$
$\quad-2\hat{n}(b_{\beta})\hat{n}(a_{\beta})\hat{n}(i_{\beta}))\hat{G}_{i_{\alpha}j_{\alpha}}^{a_{\alpha}b_{\alpha}}$
$A_{76}$ $=(\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})+4\hat{n}(b_{\alpha})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})-2\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})$
$\quad-2\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})-2\hat{n}(b_{\alpha})\hat{n}(a_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(a_{\alpha})$
$\quad-2\hat{n}(b_{\alpha})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha}))\hat{G}_{i_{\beta}j_{\beta}}^{a_{\beta}b_{\beta}}$
$A_{77}$ $=(\hat{n}(j_{\beta})\hat{n}(i_{\beta})-\hat{n}(b_{\beta})\hat{n}(i_{\beta})-\hat{n}(b_{\beta})\hat{n}(j_{\beta})+\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\alpha}}^{a_{\alpha}b_{\alpha}}$
$A_{78}$ $=(\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(a_{\alpha}))\hat{G}_{i_{\beta}j_{\beta}}^{a_{\beta}b_{\beta}}$
$A_{79}$ $=(\hat{n}(b_{\alpha})\hat{n}(j_{\beta})-\hat{n}(j_{\beta})\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})-\hat{n}(b_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{80}$ $=(\hat{n}(b_{\beta})\hat{n}(j_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(j_{\alpha})-\hat{n}(b_{\beta})\hat{n}(a_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{81}$ $=-(1-\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(i_{\alpha})\hat{n}(a_{\alpha})+\hat{n}(i_{\alpha})\hat{n}(b_{\beta})+\hat{n}(j_{\beta})\hat{n}(a_{\alpha})$
$\quad+\hat{n}(j_{\beta})\hat{n}(b_{\beta})+2\hat{n}(i_{\alpha})\hat{n}(j_{\beta})+2\hat{n}(a_{\alpha})\hat{n}(b_{\beta})-2\hat{n}(i_{\alpha})\hat{n}(a_{\alpha})\hat{n}(b_{\beta})-2\hat{n}(j_{\beta})\hat{n}(a_{\alpha})\hat{n}(b_{\beta})$
$\quad-2\hat{n}(i_{\alpha})\hat{n}(j_{\beta})\hat{n}(a_{\alpha})-2\hat{n}(i_{\alpha})\hat{n}(j_{\beta})\hat{n}(b_{\beta})+4\hat{n}(i_{\alpha})\hat{n}(j_{\beta})\hat{n}(a_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{82}$ $=-(1-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})+\hat{n}(i_{\beta})\hat{n}(a_{\beta})+\hat{n}(i_{\beta})\hat{n}(b_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(a_{\beta})$
$\quad+\hat{n}(j_{\alpha})\hat{n}(b_{\alpha})+2\hat{n}(i_{\beta})\hat{n}(j_{\alpha})+2\hat{n}(a_{\beta})\hat{n}(b_{\alpha})-2\hat{n}(i_{\beta})\hat{n}(a_{\beta})\hat{n}(b_{\alpha})-2\hat{n}(j_{\alpha})\hat{n}(a_{\beta})\hat{n}(b_{\alpha})$
$\quad-2\hat{n}(i_{\beta})\hat{n}(j_{\alpha})\hat{n}(a_{\beta})-2\hat{n}(i_{\beta})\hat{n}(j_{\alpha})\hat{n}(b_{\alpha})+4\hat{n}(i_{\beta})\hat{n}(j_{\alpha})\hat{n}(a_{\beta})\hat{n}(b_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{83}$ $=(\hat{n}(b_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(i_{\beta})-\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(i_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{84}$ $=(\hat{n}(b_{\beta})\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})\hat{n}(i_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(i_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
Commuting groups
The groups of commuting generators are,
$s$ $=\{1,2,15,16,17,18,31,32,33,34,39,40,41,42,43,44,65,66,67,68,69,70,71,72,81,82,83,84\}$ (23)
$s$ $=\{3,4,9,10,13,14,19,20,23,24,27,28,45,46,47,48,51,52,55,56,57,58,61,62,73,74,79,80\}$ (24)
$s$ $=\{5,6,7,8,11,12,21,22,25,26,29,30,35,36,37,38,49,50,53,54,59,60,63,64,75,76,77,78\}$ (25)
Algebraic structure
Let us denote $\mathfrak{s}=[\mathcal{A}_{S4T},\mathcal{A}_{S4T}]$ a derived algebra that constitute the semi-simple part, $\dim\mathfrak{s}=78$. There is a 6 dimensional center, $Z(\mathcal{A}_{S4T})$. The solvable radical coincides with the center: $\mathfrak{r}=Z(\mathcal{A}_{S4T})$, and
$\mathcal{A}_{S4T}\;\cong\;\mathfrak{s}\;\oplus\;Z(\mathcal{A}_{S4T})$
is a reductive Lie algebra (direct sum; $\mathfrak{s},Z(\mathcal{A}_{S4T})]=0$).
The algebra $\mathcal{A}_{S4T}$ is reductive with
$\dim\mathcal{A}_{S4T}=84,\qquad\dim Z(\mathcal{A}_{S4T})=6,\qquad\dim[\mathcal{A}_{S4T},\mathcal{A}_{S4T}]=78.$
It decomposes as a direct sum of ideals
$\mathcal{A}_{S4T}\;=\;Z(\mathcal{A}_{S4T})\;\oplus\;\bigoplus_{r=1}^{26}\mathfrak{s}_{r},$
where each $\mathfrak{s}_{r}$ is a $3$-dimensional simple ideal (isomorphic to $\mathfrak{sl}_{2}$ over $\mathbb{C}$), and $\mathfrak{s}_{r},\mathfrak{s}_{r^{\prime}}]=0$ for $r\neq r^{\prime}$.
Center $Z(\mathcal{A}_{S4T})$
A basis $\{Z_{1},\dots,Z_{6}\}$ of the center is:
$Z_{1}$ $=2A_{3}+2A_{19}+A_{24}+2A_{27}+2A_{48}+A_{73}$
$Z_{2}$ $=2A_{4}+2A_{20}+A_{23}+2A_{28}+2A_{46}+A_{74}$
$Z_{3}$ $=2A_{5}+2A_{21}+A_{26}+2A_{29}-2A_{38}+A_{75}$
$Z_{4}$ $=2A_{6}+2A_{22}+A_{25}+2A_{30}-2A_{37}+A_{76}$
$Z_{5}$ $=-2A_{2}+A_{32}+2A_{33}+2A_{40}-2A_{42}+A_{81}$
$Z_{6}$ $=-2A_{1}+2A_{31}+A_{34}+2A_{39}-2A_{44}+A_{82}$
Semisimple part $\mathcal{A}_{S4T},\mathcal{A}_{S4T}]$
We list the $26$ commuting simple ideals $\mathfrak{s}_{r}$ (each of dimension $3$) via generators
$\mathfrak{s}_{r}=\mathrm{span}\{B^{(r)}_{1},B^{(r)}_{2},B^{(r)}_{3}\},\qquad B^{(r)}_{m}=\sum_{k}c^{(r,m)}_{k}A_{k}.$
Ideal $\mathfrak{s}_{1}$ (dim 33).
$B^{(1)}_{1}$ $=-A_{32}-A_{34}+2A_{67}+2A_{68}+A_{81}+A_{82}$
$B^{(1)}_{2}$ $=-A_{25}-A_{26}-A_{49}-A_{50}-A_{53}-A_{54}+A_{75}+A_{76}$
$B^{(1)}_{3}$ $=-A_{23}-A_{24}+A_{51}+A_{52}+A_{55}+A_{56}+A_{73}+A_{74}$
Ideal $\mathfrak{s}_{2}$ (dim 33).
$B^{(2)}_{1}$ $=-A_{32}-A_{34}-2A_{67}-2A_{68}+A_{81}+A_{82}$
$B^{(2)}_{2}$ $=-A_{25}-A_{26}+A_{49}+A_{50}+A_{53}+A_{54}+A_{75}+A_{76}$
$B^{(2)}_{3}$ $=-A_{23}-A_{24}-A_{51}-A_{52}-A_{55}-A_{56}+A_{73}+A_{74}$
Ideal $\mathfrak{s}_{3}$ (dim 33).
$B^{(3)}_{1}$ $=A_{32}-A_{34}+2A_{67}-2A_{68}-A_{81}+A_{82}$
$B^{(3)}_{2}$ $=-A_{25}+A_{26}+A_{49}-A_{50}-A_{53}+A_{54}-A_{75}+A_{76}$
$B^{(3)}_{3}$ $=-A_{23}+A_{24}-A_{51}+A_{52}+A_{55}-A_{56}-A_{73}+A_{74}$
Ideal $\mathfrak{s}_{4}$ (dim 33).
$B^{(4)}_{1}$ $=A_{32}-A_{34}-2A_{67}+2A_{68}-A_{81}+A_{82}$
$B^{(4)}_{2}$ $=-A_{25}+A_{26}-A_{49}+A_{50}+A_{53}-A_{54}-A_{75}+A_{76}$
$B^{(4)}_{3}$ $=-A_{23}+A_{24}+A_{51}-A_{52}-A_{55}+A_{56}-A_{73}+A_{74}$
Ideal $\mathfrak{s}_{5}$ (dim 33).
$B^{(5)}_{1}$ $=-\frac{1}{2}A_{34}-A_{65}+A_{68}+\frac{1}{2}A_{82}+A_{83}$
$B^{(5)}_{2}$ $=-\frac{1}{2}A_{24}+\frac{1}{2}A_{52}+\frac{1}{2}A_{55}-A_{57}+\frac{1}{2}A_{73}+A_{79}$
$B^{(5)}_{3}$ $=\frac{1}{2}A_{26}+\frac{1}{2}A_{50}+\frac{1}{2}A_{53}+A_{59}-\frac{1}{2}A_{75}+A_{77}$
Ideal $\mathfrak{s}_{6}$ (dim 33).
$B^{(6)}_{1}$ $=-\frac{1}{2}A_{34}+A_{65}-A_{68}+\frac{1}{2}A_{82}+A_{83}$
$B^{(6)}_{2}$ $=-\frac{1}{2}A_{24}-\frac{1}{2}A_{52}-\frac{1}{2}A_{55}+A_{57}+\frac{1}{2}A_{73}+A_{79}$
$B^{(6)}_{3}$ $=\frac{1}{2}A_{26}-\frac{1}{2}A_{50}-\frac{1}{2}A_{53}-A_{59}-\frac{1}{2}A_{75}+A_{77}$
Ideal $\mathfrak{s}_{7}$ (dim 33).
$B^{(7)}_{1}$ $=-A_{31}-\frac{1}{2}A_{34}-A_{66}+A_{67}-\frac{1}{2}A_{82}+A_{83}$
$B^{(7)}_{2}$ $=A_{7}-A_{19}-\frac{1}{2}A_{24}-\frac{1}{2}A_{25}-\frac{1}{2}A_{50}+\frac{1}{2}A_{51}+\frac{1}{2}A_{53}+\frac{1}{2}A_{56}-A_{58}-A_{59}-\frac{1}{2}A_{73}+\frac{1}{2}A_{76}-A_{78}+A_{79}$
$B^{(7)}_{3}$ $=-A_{7}+\frac{1}{2}A_{25}+\frac{1}{2}A_{50}-\frac{1}{2}A_{53}+A_{59}-\frac{1}{2}A_{76}+A_{78}$
Ideal $\mathfrak{s}_{8}$ (dim 33).
$B^{(8)}_{1}$ $=-A_{31}-\frac{1}{2}A_{34}+A_{66}-A_{67}-\frac{1}{2}A_{82}+A_{83}$
$B^{(8)}_{2}$ $=-A_{19}-\frac{1}{2}A_{24}-\frac{1}{2}A_{51}-\frac{1}{2}A_{56}+A_{58}-\frac{1}{2}A_{73}+A_{79}$
$B^{(8)}_{3}$ $=A_{7}+\frac{1}{2}A_{25}-\frac{1}{2}A_{50}+\frac{1}{2}A_{53}-A_{59}-\frac{1}{2}A_{76}+A_{78}$
Ideal $\mathfrak{s}_{9}$ (dim 33).
$B^{(9)}_{1}$ $=A_{32}+A_{34}+A_{81}+A_{82}$
$B^{(9)}_{2}$ $=A_{23}-A_{24}-A_{73}+A_{74}$
$B^{(9)}_{3}$ $=-A_{49}+A_{50}-A_{53}+A_{54}$
Ideal $\mathfrak{s}_{10}$ (dim 33).
$B^{(10)}_{1}$ $=-A_{32}+A_{34}-A_{81}+A_{82}$
$B^{(10)}_{2}$ $=A_{23}+A_{24}+A_{73}+A_{74}$
$B^{(10)}_{3}$ $=-A_{49}-A_{50}+A_{53}+A_{54}$
Ideal $\mathfrak{s}_{11}$ (dim 33).
$B^{(11)}_{1}$ $=\frac{1}{2}A_{34}-A_{41}-A_{44}+A_{67}-A_{71}-\frac{1}{2}A_{82}+A_{83}$
$B^{(11)}_{2}$ $=-A_{9}-\frac{1}{2}A_{23}+\frac{1}{2}A_{52}-\frac{1}{2}A_{55}-A_{57}+\frac{1}{2}A_{74}+A_{80}$
$B^{(11)}_{3}$ $=A_{21}+\frac{1}{2}A_{26}+\frac{1}{2}A_{49}+\frac{1}{2}A_{54}+A_{60}+\frac{1}{2}A_{75}+A_{77}$
Ideal $\mathfrak{s}_{12}$ (dim 33).
$B^{(12)}_{1}$ $=\frac{1}{2}A_{34}-A_{41}-A_{44}-A_{67}+A_{71}-\frac{1}{2}A_{82}+A_{83}$
$B^{(12)}_{2}$ $=A_{9}-\frac{1}{2}A_{23}-\frac{1}{2}A_{52}+\frac{1}{2}A_{55}+A_{57}+\frac{1}{2}A_{74}+A_{80}$
$B^{(12)}_{3}$ $=A_{21}+\frac{1}{2}A_{26}-\frac{1}{2}A_{49}-\frac{1}{2}A_{54}-A_{60}+\frac{1}{2}A_{75}+A_{77}$
Ideal $\mathfrak{s}_{13}$ (dim 33).
$B^{(13)}_{1}$ $=\frac{1}{2}A_{34}+A_{39}-A_{41}-A_{44}+A_{68}-A_{72}+\frac{1}{2}A_{82}+A_{83}$
$B^{(13)}_{2}$ $=-A_{10}-A_{20}-\frac{1}{2}A_{23}+\frac{1}{2}A_{51}-\frac{1}{2}A_{56}-A_{58}-\frac{1}{2}A_{74}+A_{80}$
$B^{(13)}_{3}$ $=-A_{8}+A_{22}+\frac{1}{2}A_{25}+\frac{1}{2}A_{49}-\frac{1}{2}A_{54}+A_{60}+\frac{1}{2}A_{76}+A_{78}$
Ideal $\mathfrak{s}_{14}$ (dim 33).
$B^{(14)}_{1}$ $=\frac{1}{2}A_{34}+A_{39}-A_{41}-A_{44}-A_{68}+A_{72}+\frac{1}{2}A_{82}+A_{83}$
$B^{(14)}_{2}$ $=A_{10}-A_{20}-\frac{1}{2}A_{23}-\frac{1}{2}A_{51}+\frac{1}{2}A_{56}+A_{58}-\frac{1}{2}A_{74}+A_{80}$
$B^{(14)}_{3}$ $=A_{8}+A_{22}+\frac{1}{2}A_{25}-\frac{1}{2}A_{49}+\frac{1}{2}A_{54}-A_{60}+\frac{1}{2}A_{76}+A_{78}$
Ideal $\mathfrak{s}_{15}$ (dim 33).
$B^{(15)}_{1}$ $=-A_{16}-A_{17}+\frac{1}{2}A_{32}-A_{42}-A_{43}-A_{66}+A_{67}-A_{70}-A_{71}-\frac{1}{2}A_{81}+A_{84}$
$B^{(15)}_{2}$ $=-A_{13}+\frac{1}{2}A_{23}+\frac{1}{2}A_{26}+A_{29}-A_{35}-A_{38}+A_{46}+A_{47}-\frac{1}{2}A_{50}-\frac{1}{2}A_{51}-\frac{1}{2}A_{53}+\frac{1}{2}A_{56}-A_{61}-A_{64}-\frac{1}{2}A_{74}+\frac{1}{2}A_{75}-A_{77}+A_{80}$
$B^{(15)}_{3}$ $=-\frac{1}{2}A_{26}-A_{29}+A_{35}+A_{38}+\frac{1}{2}A_{50}+\frac{1}{2}A_{53}+A_{64}-\frac{1}{2}A_{75}+A_{77}$
Ideal $\mathfrak{s}_{16}$ (dim 33).
$B^{(16)}_{1}$ $=A_{16}+A_{17}+\frac{1}{2}A_{32}-A_{42}-A_{43}+A_{66}-A_{67}+A_{70}+A_{71}-\frac{1}{2}A_{81}+A_{84}$
$B^{(16)}_{2}$ $=A_{13}+\frac{1}{2}A_{23}+A_{46}+A_{47}+\frac{1}{2}A_{51}-\frac{1}{2}A_{56}+A_{61}-\frac{1}{2}A_{74}+A_{80}$
$B^{(16)}_{3}$ $=-\frac{1}{2}A_{26}-A_{29}+A_{35}+A_{38}-\frac{1}{2}A_{50}-\frac{1}{2}A_{53}-A_{64}-\frac{1}{2}A_{75}+A_{77}$
Ideal $\mathfrak{s}_{17}$ (dim 33).
$B^{(17)}_{1}$ $=-A_{15}-A_{18}+\frac{1}{2}A_{32}+A_{40}-A_{42}-A_{43}-A_{65}+A_{68}-A_{69}-A_{72}+\frac{1}{2}A_{81}+A_{84}$
$B^{(17)}_{2}$ $=-A_{14}+\frac{1}{2}A_{23}+A_{28}+A_{46}+A_{47}-\frac{1}{2}A_{52}+\frac{1}{2}A_{55}-A_{62}+\frac{1}{2}A_{74}+A_{80}$
$B^{(17)}_{3}$ $=-A_{12}-\frac{1}{2}A_{25}-A_{30}+A_{36}+A_{37}-\frac{1}{2}A_{50}+\frac{1}{2}A_{53}+A_{64}-\frac{1}{2}A_{76}+A_{78}$
Ideal $\mathfrak{s}_{18}$ (dim 33).
$B^{(18)}_{1}$ $=A_{15}+A_{18}+\frac{1}{2}A_{32}+A_{40}-A_{42}-A_{43}+A_{65}-A_{68}+A_{69}+A_{72}+\frac{1}{2}A_{81}+A_{84}$
$B^{(18)}_{2}$ $=A_{14}+\frac{1}{2}A_{23}+A_{28}+A_{46}+A_{47}+\frac{1}{2}A_{52}-\frac{1}{2}A_{55}+A_{62}+\frac{1}{2}A_{74}+A_{80}$
$B^{(18)}_{3}$ $=A_{12}-\frac{1}{2}A_{25}-A_{30}+A_{36}+A_{37}+\frac{1}{2}A_{50}-\frac{1}{2}A_{53}-A_{64}-\frac{1}{2}A_{76}+A_{78}$
Ideal $\mathfrak{s}_{19}$ (dim 33).
$B^{(19)}_{1}$ $=-A_{25}+A_{26}-A_{32}-A_{34}+2A_{42}+2A_{44}+A_{75}-A_{76}+A_{81}+A_{82}$
$B^{(19)}_{2}$ $=A_{25}-A_{26}-A_{75}+A_{76}$
$B^{(19)}_{3}$ $=-A_{51}+A_{52}-A_{55}+A_{56}$
Ideal $\mathfrak{s}_{20}$ (dim 33).
$B^{(20)}_{1}$ $=A_{32}-A_{34}-2A_{42}+2A_{44}-A_{81}+A_{82}$
$B^{(20)}_{2}$ $=A_{25}+A_{26}+A_{75}+A_{76}$
$B^{(20)}_{3}$ $=-A_{51}-A_{52}+A_{55}+A_{56}$
Ideal $\mathfrak{s}_{21}$ (dim 33).
$B^{(21)}_{1}$ $=-A_{25}+A_{26}+2A_{37}-2A_{38}-A_{75}+A_{76}$
$B^{(21)}_{2}$ $=-A_{23}-A_{24}-2A_{46}-2A_{48}+A_{73}+A_{74}$
$B^{(21)}_{3}$ $=-A_{15}+A_{16}-A_{65}+A_{66}-A_{67}+A_{68}-A_{69}+A_{70}$
Ideal $\mathfrak{s}_{22}$ (dim 33).
$B^{(22)}_{1}$ $=-A_{25}-A_{26}+2A_{37}+2A_{38}+A_{75}+A_{76}$
$B^{(22)}_{2}$ $=-A_{23}+A_{24}-2A_{46}+2A_{48}-A_{73}+A_{74}$
$B^{(22)}_{3}$ $=A_{15}+A_{16}+A_{65}+A_{66}-A_{67}-A_{68}+A_{69}+A_{70}$
Ideal $\mathfrak{s}_{23}$ (dim 33).
$B^{(23)}_{1}$ $=-\frac{1}{2}A_{32}+A_{68}-A_{69}+\frac{1}{2}A_{81}+A_{84}$
$B^{(23)}_{2}$ $=\frac{1}{2}A_{24}+A_{45}+A_{48}-\frac{1}{2}A_{51}-\frac{1}{2}A_{56}+A_{61}-\frac{1}{2}A_{73}+A_{79}$
$B^{(23)}_{3}$ $=-\frac{1}{2}A_{26}+A_{35}+A_{38}-\frac{1}{2}A_{49}-\frac{1}{2}A_{54}-A_{63}+\frac{1}{2}A_{75}+A_{77}$
Ideal $\mathfrak{s}_{24}$ (dim 33).
$B^{(24)}_{1}$ $=-\frac{1}{2}A_{32}-A_{68}+A_{69}+\frac{1}{2}A_{81}+A_{84}$
$B^{(24)}_{2}$ $=\frac{1}{2}A_{24}+A_{45}+A_{48}+\frac{1}{2}A_{51}+\frac{1}{2}A_{56}-A_{61}-\frac{1}{2}A_{73}+A_{79}$
$B^{(24)}_{3}$ $=-\frac{1}{2}A_{26}+A_{35}+A_{38}+\frac{1}{2}A_{49}+\frac{1}{2}A_{54}+A_{63}+\frac{1}{2}A_{75}+A_{77}$
Ideal $\mathfrak{s}_{25}$ (dim 33).
$B^{(25)}_{1}$ $=-\frac{1}{2}A_{32}-A_{33}+A_{67}-A_{70}-\frac{1}{2}A_{81}+A_{84}$
$B^{(25)}_{2}$ $=\frac{1}{2}A_{24}+A_{27}+A_{45}+A_{48}-\frac{1}{2}A_{52}-\frac{1}{2}A_{55}+A_{62}+\frac{1}{2}A_{73}+A_{79}$
$B^{(25)}_{3}$ $=A_{11}-\frac{1}{2}A_{25}+A_{36}+A_{37}+\frac{1}{2}A_{49}-\frac{1}{2}A_{54}-A_{63}+\frac{1}{2}A_{76}+A_{78}$
Ideal $\mathfrak{s}_{26}$ (dim 33).
$B^{(26)}_{1}$ $=-\frac{1}{2}A_{32}-A_{33}-A_{67}+A_{70}-\frac{1}{2}A_{81}+A_{84}$
$B^{(26)}_{2}$ $=\frac{1}{2}A_{24}+A_{27}+A_{45}+A_{48}+\frac{1}{2}A_{52}+\frac{1}{2}A_{55}-A_{62}+\frac{1}{2}A_{73}+A_{79}$
$B^{(26)}_{3}$ $=-A_{11}-\frac{1}{2}A_{25}+A_{36}+A_{37}-\frac{1}{2}A_{49}+\frac{1}{2}A_{54}+A_{63}+\frac{1}{2}A_{76}+A_{78}$